For arectanglar garden, the length is 12 ft and the width is 7 feet. Find the length of the diagonal. Round to the nearest tenth of a foot.
Answer by checkley77(12844)

This is a right triangle.
A^2+B^2=C^2
12^2+7^2=C^2
144+49=C^2
193=C^2
C=SQRT193
C=13.9 ANS. FOR THE DIAGONAL.